# SPACeR: Self-Play Anchoring with Centralized Reference Models

### v1

@@ Abstract
Developing autonomous vehicles (AVs) requires not only safety and efficiency, but also realistic, human-like behaviors that are socially aware and predictable. Achieving this requires sim agent policies that are human-like, fast, and scalable in multi-agent settings. Recent progress in imitation learning with large diffusion-based or tokenized models has shown that behaviors can be captured directly from human driving data, producing realistic policies. However, these models are computationally expensive, slow during inference, and struggle to adapt in reactive, closed-loop scenarios. In contrast, self-play reinforcement learning (RL) scales efficiently and naturally captures multi-agent interactions, but it often relies on heuristics and reward shaping, and the resulting policies can diverge from human norms. We propose human-like self-play, a framework that leverages a pretrained tokenized autoregressive motion model as a centralized reference policy to guide decentralized self-play. The reference model provides likelihood rewards and KL divergence, anchoring policies to the human driving distribution while preserving RL scalability. Evaluated on the Waymo Sim Agents Challenge, our method achieves competitive performance with imitation-learned policies while being up to 10× faster at inference and 50× smaller in parameter size than large generative models. In addition, we demonstrate in closed-loop ego planning evaluation tasks that our sim agents can effectively measure planner quality with fast and scalable traffic simulation, establishing a new paradigm for testing autonomous driving policies. Project website: https://spacer-ai.github.io/
@@ 1 INTRODUCTION
Simulation policies must satisfy two key requirements: (1) realism—behaving like humans, and (2) reactivity—responding meaningfully to other agents. These properties enable AVs to handle interactive, multi-agent scenarios. Broadly, there are two paradigms for constructing such policies: imitation learning and self-play reinforcement learning (RL).
Imitation learning starts from expert demonstrations, where recent advances include using tokenized models Wu et al. (2024); Zhang et al. (2025a) to generate realistic multi-agent behaviors or recently, diffusion models Jiang et al. (2024). The key advantage of imitation learning is that it learns directly from the human data distribution, producing realistic behavior. However, such approaches struggle in reactive settings, and often require large, computationally heavy models that limit scalability that limit scalability, as we will demonstrate in section 4.2.
@@
To address both challenges, we introduce SPACeR, which leverages a pretrained tokenized model Wu et al. (2024) to provide a reference policy distribution and a likelihood signal on the outputs of self-play policies (fig. 1). This serves as an on-policy reward provider, shaping learning toward human-like behavior while preserving the scalability of self-play RL. Specifically, we align the self-play policy’s action space with that of the tokenized model, enabling tractable likelihood estimation and KL-based distributional alignment. This design eliminates the need for ground-truth logged trajectories and instead provides probabilistic estimates that directly guide learning.
@@ 2 Related Works
Self-Play RL for autonomous driving Self-play reinforcement learning has gained traction in autonomous driving since it allows agents to learn directly from closed-loop multi-agent interaction experience Zhang et al. (2024); Kazemkhani et al. (2024); Cusumano-Towner et al. (2025). One representative work, GIGAFlow Cusumano-Towner et al. (2025) demonstrates that large-scale self-play can produce robust autonomous driving policies, showcasing the scalability and effectiveness of self-play for interactive autonomy. Nevertheless, one of the main challenges for self-play methods in real-world deployment is the divergence from human-behaviors. Policies that maximize rewards may behave unpredictably, and be unable to coordinate with humans safely. While Human-Regularized PPO Cornelisse & Vinitsky (2024) introduces a small amount of demonstration data as a regularizer to bias policies toward human-likeness, our approach leverages a large-scale tokenized model to provide a human-like reward signal during multi-agent learning, making it the first to demonstrate competitive performance with state-of-the-art imitation learning policies while preserving the scalability of self-play.
@@
Imitation-Learning Based Traffic simulation Recently, data-driven methods like imitation learning have gained popularity due to their capacity to learn from an expert data-distribution with minimal human effort. They can mainly be categroized as diffusion models and tokenized models. Diffusion Models Zhong et al. (2023); Chang et al. (2024) offer flexibitly and controllabillity to potentially generate long-tail driving scenarios. On the other hand, tokenized models such as SMARTWu et al. (2024) or CAT-KZhang et al. (2025a) have become more popular due to their high realism and capacity to simulate realistic multi-agent behavior. However, both approaches suffer from slow sampling: diffusion models rely on iterative denoising, while tokenized models depend on auto-regressive generation. This limits these imitation learning models for large-scale, closed-loop simulation. In this work, we instead utilize the prior information from pretrained imitation-learning models to guide self-play RL to generate human-like behaviors. There are several works that focus on RL fine-tuning of pretrained imitation-learning models Peng et al. (2024); Chen et al. (2025); Cao et al. (2024): for example, using Group Relative Proximal Optimization to improve realism and controllability Chen et al. (2025), leveraging human feedback for post-training alignment Cao et al. (2024), or exploiting implicit preferences from pre-training demonstrations to avoid costly human annotations Tian & Goel (2025). In contrast to this pretrain-then-finetune paradigm, our work takes an RL-first approach, where large-scale self-play serves as the foundation and imitation-learning models are incorporated as a reward provider.
@@ 3 Human-like self-play
The goal is to train a human-like driving policy $\pi_{\theta}(a|o)$ that generates behaviors resemble real human driving distribution in closed-loop. The desired policy should not only focus on tasks, such as reaching goals, avoid collisions and offroad, but produce behaviors that are predictable, and realistic.
@@ 3.1 Problem Formulation
We formulate this as a partially observable multi-sequential decision-making problem. At each timestep $t$, the global simulator state $s_{t}\in\mathcal{S}$ encodes the static road graph and the dynamic states of all agents. Each agent $i$ receives only a partial observation $o_{t}^{i}=\mathcal{O}(s_{t},i)$ within its local field-of-view. The agent then chooses an action $a_{t}^{i}\in\mathcal{A}$ (e.g., acceleration, steering, or maneuver) according to its policy $\pi(a_{t}^{i}|o_{t}^{i})$, and the environment transitions according to $s_{t+1}\sim T(s_{t},a_{t}^{1},\dots,a_{t}^{N})$.
@@
$\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{PPO}}(\theta;A[r])-\beta\,D_{\mathrm{KL}}\!\big(\pi_{\mathrm{ref}}(\cdot\mid s_{t})\,\|\,\pi_{\theta}(\cdot\mid o_{t})\big),$ (2)
@@
  • Distributional alignment: The KL divergence $D_{\mathrm{KL}}(\pi_{\mathrm{ref}}(\cdot\mid s_{t})\,\|\,\pi_{\theta}(\cdot\mid o_{t}))$ acts as a dense signal at every timestep, directly encouraging the policy to align its action distribution with the human driving distribution captured by the reference model.
@@ 3.2 Pretrained Reference Tokenized Model
To incorporate human-likeness into self-play, we introduce a pretrained reference policy $\pi_{\text{ref}}$, trained on real-world human driving trajectories, to serve as a proxy for the human driving distribution. Unlike task-based rewards, which only encourage efficiency and safety, $\pi_{\text{ref}}$ provides a distributional signal that anchors policies toward realistic behavior. In principle, $\pi_{\text{ref}}$ can be any data-driven generative motion model, such as a diffusion mode or a tokenized model. In this work, we focus on tokenized models (e.g., SMART Wu et al. (2024), CAT-K Zhang et al. (2025a)), since they provide tractable likelihood estimates, and capability to provide distributional signal.
@@
$D_{\mathrm{KL}}\!\big(\pi_{\mathrm{ref}}(\cdot\mid s_{t})\,\|\,\pi(\cdot\mid o_{t})\big)=\sum_{a\in\mathcal{A}}\pi_{\mathrm{ref}}(a\mid s_{t})\,\log\frac{\pi_{\mathrm{ref}}(a\mid s_{t})}{\pi(a\mid o_{t})}.$
(3) Privileged information and comparison to log-replay data. Note that in Eq. 2, the reference tokenized model is centralized, observing the full state of all agents rather than local views. This setup resembles privileged information in a teacher–student framework in Robotics and Computer Vision He et al. (2024); Caron et al. (2021). Compared with log-replay trajectories, which are restricted to recorded behaviors, the reference model can generalize beyond the dataset and provide broader distributional guidance.
Reference Models vs. WOSAC Metrics WOSAC metrics Wu et al. (2024) estimate per-feature likelihoods (kinematic, interactive, map-based) based on ground-truth future trajectories. In contrast, tokenized reference models directly define a distribution over entire action sequences, offering two advantages: (i) evaluation does not require logged future trajectories, and (ii) the model provides a principled sequence-level distribution rather than feature-wise likelihoods tied to manual design choices.
@@ 3.3 Practical Considerations of building human-like self-play
Goal-reaching reward and post-goal behavior In previous works Kazemkhani et al. (2024); Cornelisse et al. (2025), agents are typically rewarded only upon reaching their goal, and are removed from the scenario once the goal is reached. This formulation has two main drawbacks: (i) agents are incentivized to accelerate unnaturally in order to reach the goal as quickly as possible, and (ii) the number of active agents decreases after goals are reached, making the scenario progressively easier. Qualitatively, with this original formulation, if the agents are not disappeared, agents would mostly stop once reached the goal, making the scenarios unrealistic.
To address this issue, we adopt goal-dropout during training: agents are trained both with and without goal conditioning, while still receiving a goal-reaching reward only at the end of the episode. This forces agents to implicitly predict their own goals rather than relying on privileged information. Moreover, in our experiments we show that with the reference tokenized model, the explicit goal-reaching reward can be dropped entirely.
@@ 4.1 Implementation Details
We conduct all experiments in GPUDrive Kazemkhani et al. (2024), a GPU-accelerated, data-driven simulator built on WOMD Ettinger et al. (2021). Each WOMD scenario spans 9 seconds; following the setup of Montali et al. (2023), we initialize at 1 second and simulate the remaining 8 seconds. We train on 10k scenarios. In this work, we only control vehicles, while pedestrians and cyclists follow their logged trajectories.
@@
Action Space. We adopt a tokenized trajectory action space following Wu et al. (2024); Philion et al. (2024). Using the K-disk algorithm, we cluster short-horizon trajectories in Cartesian space into $K=200$ discrete tokens. Each token represents a 0.1-second step with a horizon length of 2, corresponding to a 5 Hz action frequency. This setup balances two considerations: providing sufficiently fine-grained distributional signals from the SMART reference model while keeping memory usage manageable. Since actions are defined directly in Cartesian space, no explicit dynamics model is assumed; the simulator advances the scenario according to the selected token.
@@
$r^{\text{task}}(o_{t},a_{t})=w_{\text{goal}}\,\mathbb{I}[\text{goal achieved}]-w_{\text{collided}}\,\mathbb{I}[\text{collision}]-w_{\text{offroad}}\,\mathbb{I}[\text{offroad}]+w_{\text{humanlike}}\,r^{\text{humanlike}}(s_{t},a_{t})$
where $\mathbb{I}[\cdot]\in\{0,1\}$. By default, we set $w_{\text{collided}}=w_{\text{offroad}}=0.75$. The weights $w_{\text{goal}}$ and $w_{\text{humanlike}}$ are varied in the following ablation experiments to study the trade-off between task completion, safety, and human-likeness. Note that other than the reward, we also directly compare the KL divergence between trained policy and referenced policy, where we use Cat-K Zhang et al. (2025a) as the reference model.
Model Architecture and Training Details: We use a late-fusion feedforward model Kazemkhani et al. (2024); Cornelisse et al. (2025), where ego, partner, and road-graph features are each embedded with a two-layer MLP and then concatenated. The fused representation is fed into an actor head (action logits) and a critic head (value). During training, we control up to 64 agents with a shared decentralized policy $\pi_{\theta}$, where each agent samples actions based on its local observation. We optimize with Proximal Policy Optimization (PPO) Schulman et al. (2017), with hyperparameters provided in section A.1. Training is performed on a single NVIDIA A100 GPU for 1 billion environment steps. During training, all VRUs (pedestrian, cyclists) are not controlled and follow the logs.
4.2 Humanlike self-play Evaluation on WOSAC
@@
HR-PPO Cornelisse & Vinitsky (2024). PPO regularized toward a behavioral cloning (BC) reference using KL divergence only (no likelihood term). For fairness, the BC model shares the same backbone as $\pi$, is decentralized, and conditions only on local observations; it is trained on the WOMD training split.
Imitation learning baselines. We evaluate two tokenized closed-loop models—SMART Wu et al. (2024) and CAT-K Zhang et al. (2025b). Both use the same backbone and action vocabulary (K=200K\!=\!200) and run at 5 Hz. SMART is first trained by behavior cloning on WOMD and then fine-tuned following the CAT-K closed-loop fine-tuning protocal; Implementation details and hyperparameters are provided in section A.1.
Discussion of Results. In table 1, anchoring self-play with a pretrained reference model substantially improves realism across all Sim Agent metrics. By contrast, HR-PPO reduces minADE but yields only a modest realism gain, whereas our method achieves clear improvements in kinematics, underscoring the value of alignment with a strong reference model.
@@
Variant Composite $\uparrow$ minADE $\downarrow$
No KL and LLH 0.70 14.43
Goal + LLH (no KL) 0.69 21.05
Goal + KL 0.73 4.08
KL + $r_{\text{infraction}}$ (no goal) 0.74 4.73
KL + $r_{\text{infraction}}$ + LLH (no goal) 0.74 4.68
Table 2: Ablation on WOSAC (validation). $r_{\text{infraction}}$: off-road/collision penalties; LLH: log-likelihood reward. KL is the dominant factor for realism.
In the ablation study (table 2), KL alignment contributes most to realism. While log-likelihood rewards are popular Escontrela et al. (2023), their impact on realism is modest, likely because the real-world driving distribution is highly multi-modal, so maximizing likelihood alone does not yield a stable or sufficiently discriminative signal. A second insight is that, once the policy is anchored to the reference distribution, the explicit goal reward can be removed, which further improves realism.

### v2

@@ Abstract
Developing autonomous vehicles (AVs) requires not only safety and efficiency, but also realistic, human-like behaviors that are socially aware and predictable. Achieving this requires sim agent policies that are human-like, fast, and scalable in multi-agent settings. Recent progress in imitation learning with large diffusion-based or tokenized models has shown that behaviors can be captured directly from human driving data, producing realistic policies. However, these models are computationally expensive, slow during inference, and struggle to adapt in reactive, closed-loop scenarios. In contrast, self-play reinforcement learning (RL) scales efficiently and naturally captures multi-agent interactions, but it often relies on heuristics and reward shaping, and the resulting policies can diverge from human norms. We propose human-like self-play, a framework that leverages a pretrained tokenized autoregressive motion model as a centralized reference policy to guide decentralized self-play. The reference model provides likelihood rewards and KL divergence, anchoring policies to the human driving distribution while preserving RL scalability. Evaluated on the Waymo Sim Agents Challenge, our method achieves competitive performance with imitation-learned policies while being up to 10× faster at inference and 50× smaller in parameter size than large generative models. In addition, we demonstrate in closed-loop ego planning evaluation tasks that our sim agents can effectively measure planner quality with fast and scalable traffic simulation, establishing a new paradigm for testing autonomous driving policies.
@@ 1 INTRODUCTION
Simulation policies must satisfy two key requirements: realism and reactivity. Realism refers to human-like behavior, while reactivity captures meaningful responses to other agents. These properties enable AVs to handle interactive, multi-agent scenarios. Broadly, there are two paradigms for constructing such policies: imitation learning and self-play reinforcement learning (RL).
Imitation learning starts from expert demonstrations, with recent advances including tokenized models Philion et al. (2024); Wu et al. (2024); Zhang et al. (2025) and diffusion models Zhong et al. (2023); Jiang et al. (2024) to generate realistic multi-agent behaviors. The key advantage of imitation learning is that it learns directly from the human data distribution, producing realistic behavior. However, such approaches struggle in reactive settings and often require computationally heavy models such as Transformer-based architectures that limit scalability, as we demonstrate in section 4.3.
@@
To address both challenges, we introduce SPACeR, which leverages a pretrained tokenized model Wu et al. (2024) to provide a reference policy distribution and a likelihood signal over the outputs of self-play policies (fig. 1). This serves as an on-policy reward provider, shaping learning toward human-like behavior while preserving the scalability of self-play RL. Specifically, we align the self-play policy’s action space with that of the tokenized model, enabling tractable likelihood estimation and KL-based distributional alignment. This design eliminates the need for ground-truth logged trajectories and instead provides probabilistic estimates that directly guide learning.
@@ 2 Related Works
Self-play reinforcement learning has been studied extensively in multi-agent reinforcement learning (MARL) as a way for agents to learn by interacting with copies of themselves Carroll et al. (2019); Yu et al. (2022); Chen et al. (2024) and has recently been applied to autonomous driving Zhang et al. (2024); Kazemkhani et al. (2024); Cusumano-Towner et al. (2025). One representative work, GIGAFlow Cusumano-Towner et al. (2025) demonstrates that large-scale self-play can produce robust autonomous driving policies, showcasing the scalability and effectiveness of self-play for autonomy. Nevertheless, one of the main challenges for self-play methods in real-world deployment is the divergence from human behaviors. Policies that maximize rewards may behave unpredictably, and be unable to coordinate with humans safely. While Human-Regularized PPO Cornelisse and Vinitsky (2024) introduces a small amount of demonstration data as a regularizer to bias policies toward human-likeness, our approach leverages a large-scale tokenized model to provide a human-like reward signal during multi-agent learning, making it the first to demonstrate competitive performance with state-of-the-art imitation learning policies while preserving the scalability of self-play.
Imitation-Learning Based Traffic simulation
@@
Recently, data-driven methods like imitation learning have gained popularity due to their capacity to learn from an expert data-distribution with minimal human effort. Early works have explored generative imitation-learning approaches for modeling human traffic behaviors Kuefler et al. (2017); Bhattacharyya et al. (2020). Recent works can mainly be categorized as diffusion models and tokenized models. Diffusion Models such as  Zhong et al. (2023); Chang et al. (2024; 2025); Jiang et al. (2024) offer flexibility and controllability to potentially generate long-tail driving scenarios. On the other hand, tokenized models such as SMART Wu et al. (2024) or CAT-K Zhang et al. (2025) have become more popular due to their high realism and capacity to simulate realistic multi-agent behavior. However, both approaches are computationally expensive at inference time: diffusion models require multiple denoising steps, and tokenized models generate actions sequentially. This limits scalability in large-scale closed-loop simulation. In this work, we instead leverage pretrained imitation-learning models to guide self-play RL toward human-like behaviors. There are several works that focus on RL fine-tuning of pretrained imitation-learning models Peng et al. (2024); Chen et al. (2025); Cao et al. (2024): for example, using Group Relative Proximal Optimization to improve realism and controllability Chen et al. (2025), leveraging human feedback for post-training alignment Cao et al. (2024), or exploiting implicit preferences from pre-training demonstrations to avoid costly human annotations Tian and Goel (2025). In contrast to this pretrain-then-finetune paradigm, we take an RL-first approach, where self-play serves as the foundation and imitation-learning models are incorporated as a reward provider.
@@ 3 Human-like self-play
We aim to learn a human-like driving policy $\pi_{\theta}(a|o)$ whose behaviors match the real human driving distribution in closed-loop. The desired policy should reach goals, avoid collisions, and stay on-road, while producing predictable and realistic behaviors.
@@ 3.1 Problem Formulation
We formulate this as a partially observable multi-sequential decision-making problem. At each timestep $t$, the global simulator state $s_{t}\in\mathcal{S}$ encodes the static road graph and the dynamic states of all agents. Each agent $i$ receives only a partial observation $o_{t}^{i}=\mathcal{O}(s_{t},i)$ within its local field-of-view. The agent then chooses an action $a_{t}^{i}\in\mathcal{A}$ (e.g., acceleration, steering, or maneuver) according to its policy $\pi_{\theta}(a_{t}^{i}|o_{t}^{i})$, and the environment transitions according to $s_{t+1}\sim T(s_{t},a_{t}^{1},\dots,a_{t}^{N})$.
@@
$\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{PPO}}(\theta;A[r])-\beta\,D_{\mathrm{KL}}\!\big(\pi_{\theta}(\cdot\mid o_{t})\,\|\,\pi_{\mathrm{ref}}(\cdot\mid s_{t})\big),$ (2)
@@
  • Distributional alignment: The KL divergence $D_{\mathrm{KL}}(\pi_{\theta}(\cdot\mid o_{t})\,\|\,\pi_{\mathrm{ref}}(\cdot\mid s_{t}))$ acts as a dense signal at every timestep, directly encouraging the policy to align its action distribution with the human driving distribution captured by the reference model. For notational simplicity, we omit the agent index $i$; the KL divergence is computed independently for each agent.
@@ 3.2 Pretrained Reference Tokenized Model
To incorporate human-likeness into self-play, we introduce a pretrained reference policy $\pi_{\text{ref}}$, trained on real-world human driving trajectories, to serve as a proxy for the human driving distribution. Unlike task-based rewards, which only encourage efficiency and safety, $\pi_{\text{ref}}$ provides a distributional signal that anchors policies toward realistic behavior. In principle, $\pi_{\text{ref}}$ can be any data-driven generative motion model, such as a diffusion model or a tokenized model. In this work, we focus on tokenized models (e.g., SMART Wu et al. (2024), CAT-K Zhang et al. (2025)), since they provide tractable likelihoods estimates suitable for distributional alignment.
@@
$D_{\mathrm{KL}}\!\big(\pi_{\theta}(\cdot\mid o_{t})\,\|\,\pi_{\mathrm{ref}}(\cdot\mid s_{t})\big)=\sum_{a\in\mathcal{A}}\pi_{\theta}(a\mid o_{t})\,\log\frac{\pi_{\theta}(a\mid o_{t})}{\pi_{\mathrm{ref}}(a\mid s_{t})}.$ (5)
(3) Privileged information and comparison to log-replay data. Note that in Eq. 2, the reference tokenized model is centralized, observing the full state of all agents rather than local views. This setup resembles privileged information in a teacher–student framework in Robotics and Computer Vision He et al. (2024); Caron et al. (2021). Unlike log replay, the reference model generalizes beyond recorded trajectories and provides guidance in unseen self-play states.
Reference Models vs. WOSAC Metrics. WOSAC metrics Wu et al. (2024) estimate per-feature likelihoods (kinematic, interactive, and map-based) based on ground-truth future trajectories. In contrast, tokenized reference models directly define a distribution over entire action sequences, offering two advantages: (i) evaluation does not require logged future trajectories, and (ii) the model provides a principled sequence-level distribution rather than feature-wise likelihoods tied to manual design metrics.
@@ 3.3 Practical Considerations of building human-like self-play
Goal-reaching reward and post-goal behavior In previous works Kazemkhani et al. (2024); Cornelisse et al. (2025), agents are typically rewarded only upon reaching their goal, and are removed from the scenario once the goal is reached. This formulation has two main drawbacks: (i) agents are incentivized to accelerate unnaturally in order to reach the goal as quickly as possible, and (ii) the number of active agents decreases after goals are reached, making the scenario progressively easier. Qualitatively, with this original formulation, if the agents are not disappeared, agents would mostly stop once reached the goal, making the scenarios unrealistic. To address this issue, we adopt goal-dropout: agents are trained with and without goal conditioning while receiving a terminal goal reward. This reduces reliance on explicit goal inputs. We further show that, when anchored to a reference model, the explicit goal reward can be removed without performance loss.
@@ 4.1 Implementation Details
We conduct all experiments in GPUDrive Kazemkhani et al. (2024), a GPU-accelerated, data-driven simulator built on WOMD Ettinger et al. (2021). Each WOMD scenario spans 9 seconds; following the setup of Montali et al. (2023), we initialize at 1 second and simulate the remaining 8 seconds. We train on 10k scenarios. In the main vehicle experiments (table 1), we control only vehicles, while pedestrians and cyclists follow their logged trajectories. For the VRU experiments (table 2), we train SPACeR by controlling all agent types, , but report metrics only on pedestrian/cyclist targets.
@@
Action Space. We adopt a tokenized trajectory action space following Philion et al. (2024); Wu et al. (2024). Using the K-disk algorithm, we cluster short-horizon trajectories in Cartesian space into $K=200$ discrete tokens. Each token represents a 0.1-second step with a horizon length of 2, corresponding to a 5 Hz action frequency. This setup balances two considerations: providing sufficiently fine-grained distributional signals from the SMART reference model while keeping memory usage manageable. Since actions are defined directly in Cartesian space, no explicit dynamics model is assumed; the simulator advances the scenario according to the selected token.
@@
$r^{\text{task}}(s_{t},a_{t})\;=\;w_{\text{goal}}\,\mathbb{I}[\text{Goal achieved}]-w_{\text{collided}}\,\mathbb{I}[\text{Collided}]-w_{\text{offroad}}\,\mathbb{I}[\text{Offroad}]+w_{\text{humanlike}}\,r^{\text{humanlike}}(s_{t},a_{t}),$
where $\mathbb{I}[\cdot]\in\{0,1\}$. By default, we set $w_{\text{collided}}=w_{\text{offroad}}=0.75$. The weights $w_{\text{goal}}$ and $w_{\text{humanlike}}$ are varied in the following ablation experiments to study the trade-off between task completion, safety, and human-likeness. Note that other than the reward, we also directly compare the KL divergence between trained policy and referenced policy, where we use CAT-K Zhang et al. (2025) as the reference model.
Model Architecture and Training Details: We use a late-fusion feedforward model Kazemkhani et al. (2024); Cornelisse et al. (2025), where ego, partner, and road-graph features are each embedded with a two-layer MLP and then concatenated. The fused representation is fed into an actor head and a critic head. During training, we control up to 64 agents with a shared decentralized policy $\pi_{\theta}$, where each agent samples actions based on its local observation. We optimize with Proximal Policy Optimization (PPO) Schulman et al. (2017), with hyperparameters provided in section A.3. Training is performed on a single NVIDIA A100 GPU for 1 billion environment steps. During training, all VRUs (pedestrians and cyclists) follow the logged trajectories except for the VRU experiments in Table 2, where we train SPACeR for all agent types, and calculate metrics on VRUs.
@@ 4.2 Humanlike self-play Evaluation on WOSAC
HR-PPO Cornelisse and Vinitsky (2024). PPO regularized toward a behavioral cloning (BC) reference using KL divergence only (no likelihood term). For fairness, the BC model shares the same backbone as $\pi_{\theta}$, is decentralized, and conditions only on local observations; it is trained on the WOMD training split.
Imitation learning baselines. We evaluate two tokenized closed-loop models—SMART Wu et al. (2024) and CAT-K Zhang et al. (2025). Both use the same backbone and action vocabulary (K=200K\!=\!200) and run at 5 Hz. SMART is first trained by behavior cloning on WOMD and then fine-tuned following the CAT-K closed-loop fine-tuning protocol; Implementation details and hyperparameters are provided in section A.3.
Figure 2: Reference Model Quality Effect on SPACeR Realism. SPACeR maintains high composite realism ( 0.73) even when anchored to a reference model of only 0.3 M parameters with realism 0.64.
Discussion of Main Results. In table 1, anchoring self-play with a pretrained reference model substantially improves realism across all Sim Agent metrics. By contrast, HR-PPO reduces minADE but yields only a modest realism gain, whereas our method achieves clear improvements in kinematics, underscoring the value of alignment with a strong reference model. Relative to imitation-learning methods, our approach attains lower collision and off-road rates and even surpasses SMART in composite realism.
@@
Reference Model Quality. We further study how reference model quality affects SPACeR performance by training SMART models of three sizes: 0.3M, 1M, and 3M parameters. Each model is also fine-tuned with CAT-K, producing a total of six reference models (fig. 2). SPACeR policy still reaches 0.732 realism even with the smallest underperforming reference model (0.636). We observe a similarly stable performance when varying the token vocabulary size with 3M params (100, 200, 400; different markers in fig. 2): larger vocabularies increase reference model’s realism, but the resulting SPACeR policies all remain clustered around 0.73. This indicates that the reference model serves mainly as a soft prior that guides humanlike behavior, rather than a target for direct imitation.
Because SPACeR agents learn through closed-loop interaction, they can refine their behavior beyond the limitations of the reference model. In addition, using a reference model that has itself undergone closed-loop fine-tuning (CAT-K) provides stronger behavioral signals, which further improves the resulting SPACeR agents performance.
Method Composite $\uparrow$ Kinematic $\uparrow$ Interactive $\uparrow$ Map $\uparrow$ minADE $\downarrow$
PPO 0.648 0.242 0.683 0.835 7.712
HR-PPO 0.668 0.285 0.700 0.847 7.014
SPACeR (Ours) 0.729 0.413 0.762 0.866 2.066
Table 2: VRU realism metrics on WOSAC (pedestrians and cyclists). SPACeR outperforms PPO and HR-PPO by a large margin, achieving substantial gains across all realism metrics and minADE.
Evaluation on Pedestrian and Cyclist Behavior. We evaluate SPACeR on the WOSAC validation set restricted to pedestrians and cyclists to isolate VRU behavior (table 2). Although absolute composite realism scores are 0.1 lower than for vehicles, SPACeR still improves performance by a large margin over PPO and HR-PPO across all metrics, including Composite, Kinematic, Interactive, and Map realism, as well as minADE. This demonstrates that the anchoring mechanism remains effective even under the higher stochasticity and variability of VRU motion. Additional design choices and component ablations for VRU simulation are provided in Appendix section A.1.
Effect of Anchoring Parameters on SPACeR Performance. We investigate the effect of the anchoring parameters, namely the likelihood weight $\alpha$ and the KL alignment weight $\beta$, by activating each term independently. Top-1 refers to evaluating agents by always taking the single most probable action, whereas top-5 evaluates realism under stochastic sampling from the top five most probable actions. Since SimAgent measures distributional realism over 32 joint rollouts, maintaining diversity through top-$k$ sampling generally yields higher composite scores Montali et al. (2023).
Figure 3: Anchoring parameter ablation. Likelihood-only improves top-1, while KL alignment improves both top-1 and top-5 realism.
The likelihood and KL alignment terms play different roles in shaping policy behavior. Optimizing the likelihood term alone increases log likelihood but reduces diversity, as reflected by a drop in entropy in the training curves (Appendix fig. A1). This produces small improvements in top-1 composite realism but harms top-5 performance. In contrast, the KL alignment term raises both likelihood and KL consistency while preserving entropy, which improves both top-1 and top-5 realism. The training curves in Appendix fig. A1 illustrate how log likelihood, KL divergence, and entropy evolve under each parameter setting.
In the ablation study (table A2), KL alignment contributes most to realism. While log-likelihood rewards are popular Escontrela et al. (2023), their impact on realism is modest, likely because the real-world driving distribution is highly multi-modal, so maximizing likelihood alone does not yield a stable or sufficiently discriminative signal. A second insight is that once the policy is anchored to the reference distribution, the explicit goal-reaching reward can be removed, further improving realism.